OpenStudy (anonymous):

A set of data points has an equation of = -1.5 + 2.5x. What is the residual for the point (5,12)? A. 11 B. -1 C. 1 D. 12

OpenStudy (anonymous):

difference between the actual \(y\) value and the predicted \(y\) value i.e. plug in \(5\) in to \(y=-1.5 + 2.5x\) for \(x\) and find what \(y\) you get

OpenStudy (anonymous):

i get \(y=-1.5+2.5\times 5=11\)

OpenStudy (anonymous):

i got 14

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (anonymous):

but instead of \(11\) you have \(12\) which lies one unit above \(11\)

OpenStudy (anonymous):

i don't know, i used a calculator

OpenStudy (anonymous):

but we can do the arithmetic if you like

OpenStudy (anonymous):

\[-1.5\times 2.5\times 5=-1.5+12.5=11\]

OpenStudy (anonymous):

in any case since \(12\) is one unit higher, the "residual' is \(1\)

OpenStudy (anonymous):

so i just plug in the y value in for x ?

OpenStudy (anonymous):

oh you added!!

OpenStudy (anonymous):

so the x value goes in the x spot right? and it supposed to equal the y value?

OpenStudy (anonymous):

steps are 1) plug in \(5\) for \(x\) in the formula 2) compare the number you get with the \(y\) coordinate of \((5,12)\) i.e. compare it to \(12\)

OpenStudy (anonymous):

got it ! thank you.

OpenStudy (anonymous):

since \(12\) is one unit above the \(11\) you get from the formula, the "residual" is one

OpenStudy (anonymous):

and for this one i would bet an even $10 that it is right yw

OpenStudy (anonymous):

you were right :)

OpenStudy (anonymous):

does the y mean anything special? in the equation

OpenStudy (anonymous):

on the link ^^

OpenStudy (anonymous):

you know i don't actually know squat about data analysis, i just google "residual" to get your last answer for this one, i would just replace \(x\) by \(15\) and see what you get i can't imagine it can be anything else

OpenStudy (anonymous):

alright thank you